A lead atom has a mass of 3.4 x 10 g neg 22 power. How many lead atoms are in a cube of lead that has a volume of 7.00cm to the 3rd power if the density of lead is 11.3g/cm to the 3rd power
1
Expert's answer
2017-01-23T06:43:08-0500
1 . ᵽ = m/V, m = ᵽ * V
m = 7*11,3 = 79,1g
2 . N = m/ma = 79,1/34*10⁻²² = 23,265*10²²